# Running in circles: practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

### v2

@@ 3.3 Application to single-cell RNA-seq data analysis
Single-cell RNA-seq (scRNA-seq) data analysis pipelines often involve an initial cluster analysis followed by differential analysis to identify marker genes and annotate cell populations based on gene expression. Given their overdispersed count nature, the negative binomial distribution is favored over the Gaussian distribution to model scRNA-seq data. Unfortunately, the negative binomial distribution also raises challenges for data thinning. The overdispersion parameter $\theta$ plays a role similar to the variance parameter in the Gaussian distribution, as it is considered known. Therefore, the quality of its estimation is directly impacts $\operatorname{Cov}\left(X^{(1)},X^{(2)}\right)=\tau(1-\tau)\frac{\mu^{2}}{\theta}\left(1-\frac{\theta+1}{\widehat{\theta}+1}\right)$, where $\widehat{\theta}$ is an estimated of $\theta$ (Neufeld et al., 2023a, ). Also, for negative binomial mixtures, data thinning decomposition is once again feasible only at the component level. As the overdispersion parameter is component-specific (Li et al.,, 2018), providing an estimator that ensures independence is here also a harduous and circular task, given that the components themselves are again unknown and thus require estimation through data thinning.
@@
We further illustrate the necessity of applying intra-component data thinning (with the associated intra-components overdispersion) to ensure Type I error control in the post-clustering inference setting with numerical simulations. We generated $n=100$ observations from a two-component negative binomial mixture: $0.5\text{NegBin}(\mu_{1},\theta_{1})+0.5\text{NegBin}(\mu_{2},\theta_{2})$ with component parameters $(\mu_{1},\theta_{1})=(5,5)$ and $(\mu_{2},\theta_{2})=(60,40)$. We conducted similar post-clustering inference as in Figure 2. In Figure 5A, applying the $k-$means with $K=3$ clusters reveals that the first mixture component is erroneously split into 2 clusters (labeled $C_{1}$ and $C_{3}$). We then assessed the Type I error rate associated with the Wilcoxon test between these incorrect clusters when applying data thinning with various overdispersion estimators. First, we applied intra-component data thinning using oracle estimates $\widetilde{\theta}_{k}$, $k=1,2$, representing true intra-component overdispersion parameters (infeasible in real-life application where the true cluster structure is unknown). We compared those results with two alternatives that are feasible in practice: i) applying intra-cluster data thinning based on the $k-$means results from Figure 5A with their associated $\widehat{\theta}_{\hat{k}}$, $\hat{k}=1,2,3$, and global data thinning with its associated $\widehat{\theta}$. Of note, all the overdispersion estimations were performed using Maximum Likelihood. Figure 5B presents the QQ-plot against the Uniform distribution of the associated Wilcoxon p-values across $1,\!000$ simulations. Similarly to the Gaussian setting, achieving uniformly distributed p-values is only possible through intra-component data thinning using their oracle overdispersion estimates $\widetilde{\theta}_{k}$. All other data thinning approaches are performed with biased estimators of the overdispersion compromising the independence between $X^{(1)}$ and $X^{(2)}$, and therefore leading to failure in the control of the post-clustering Type I error rate.
@@
Figure 5: Challenges in estimating overdispersion, a gene-specific parameter, for negative binomial data thinning. Panel A showcases erroneous clustering results on a simulated dataset, while Panel B presents a QQ-plot against the Uniform distribution, displaying Wilcoxon p-values from various data thinning approaches over $1,\!000$ simulations when testing between the two erroneous clusters. In Panel C, the estimated overdispersions of genes within different cell populations (macrophages, monocytes, granulocytes, and CD4) are plotted against those in neutrophils, along with their associated Root Mean Square Error (RMSE) for comparison.
@@
Extending our investigation from simulated scenarios to real-life applications, we used a single-cell RNA-seq dataset from the Tabula Sapiens Consortium (Consortium* et al.,, 2022) to delve into the practical challenges of estimating overdispersion. Our analysis focused on five distinct cell populations: $2,\!560$ neutrophils, $105$ macrophages, $386$ monocytes, $454$ granulocytes, and $833$ CD4 T cells – all collected from a single donor. In this controlled setting, where cell types were known, we succeeded in estimating the overdispersion of $8,\!333$ genes for each cell type using the variance stabilizing transformation implemented in the sctransform package (Choudhary and Satija,, 2022). Figure 5C illustrates the comparison of gene overdispersion when estimated solely in neutrophils versus the overdispersion estimated for the same genes in the other four cellular populations. Root Mean Squared Error (RMSE) values were computed to quantify the agreement between estimations. Our findings reveal that overdispersion is specific to each cell population, as evidenced by a notable deviation from the diagonal and relatively high RMSE values. This underscores the challenge of accurately estimating this parameter without prior knowledge of the true mixture underlying the data. Combined with our simulation studies, these results demonstrate how Type I error can easily be inflated in real-life applications of data thining for scRNA-seq data analysis, particularly due to the difficulty in providing an unbiased estimate of gene overdispersion.
@@ 4 Discussion
We highlight here the practical limitations inherent in data fission and its extension, data thinning, for post-clustering inference challenges. A crucial issue is the assumption of a homogeneous data distribution, which implies an absence of true clusters in the data. To address this limitation and adapt to scenarios with true classes, a shift towards mixture models becomes imperative. However, these models lack a predefined decomposition through data fission or thinning.
@@
We have proposed an intra-component decomposition for data fission and data thining, and demonstrated its theoretical validity. It relies on a priori knowledge of the mixture components scale parameters, such as variances in the Gaussian distribution or the overdispersion in the negative binomial distribution. However in real-life applications, these parameters are unknown. Adequately estimating these parameters becomes intricate in the presence of true clusters, given their component-specific nature; meanwhile the quality of the estimation of those parameters is directly linked with the covariance between the new random variables, $X^{(1)}$ and $X^{(2)}$, decomposing the original data. Only unbiased estimation of these parameters ensures the independence between $X^{(1)}$ and $X^{(2)}$. That independence is paramount for post-clustering inference to adequately control the Type I error rate.
@@
In the Gaussian framework, we theoretically quantify the relationship between the relative bias in variance estimation and the associated Type I error in post-clustering $t$-tests. In practice, our simulation results suggest that a small relative bias can be acceptable while still achieving effective Type I error control. These first results pave the way for defining a principled approach to tuning the hyperparmater $\tau$ in data fission and data thining for post-clustering inference to optimize statistical power.
As a solution to avoid the need for prior knowledge of component-specific variance parameters in the Gaussian mixture setting, we propose a heteroscedastic model with individual variances, that can be replaced by a plug-in estimate (such as a non-parametric estimator of the local variance). This approach aligns more closely with the distributional assumptions made by data fission and thinning. However, the performance of our non-parametric approach relies heavily on the choice of its bandwidth. The best bandwidth would be the one capturing only the observations originating from the same component, but it would again require knowledge of the true mixture components in spite of their estimation being part of the first clustering step of the method. Consequently, while this heteroscedastic model better fits the parametric assumptions of data fission and thinning, leveraging an accurate enough plug-in estimator remains challenging. This underscores the difficulty of adapting these methods to mixture distributions. We show that when the signal to noise ratio is extremely favorable for the clustering, with very well-separated components, this approach unlocks the use of data-fission for post-clustering differential testing of clustered data. In practice though, one can question the need for data fission in such cases as the uncertainty regarding the clustering is actually very small when $\delta/\sigma\geq 10$. We have also investigated iteratively fissioning the data and updating the variance plug-in estimate, but this heuristic strategy requires even larger separation of the components to effectively work.
@@
Finally, we demonstrate that the results derived from the Gaussian distribution context are readily applicable to the negative binomial distribution commonly employed in modeling RNA-seq data. Specifically, we illustrate on real data that overdispersion is also component-specific. Therefore, without knowing the true cluster structure in the data, data thinning cannot ensure the needed independence between clustering and differential testing.
@@
In practice, the application of data fission or data thinning for post-clustering inference appears to be akin to a self-referential loop generating circular reasoning. While introduced as a solution for addressing challenges in post-clustering inference, all strategies that could theoretically ensure independence between the two stages of the analysis ultimately rely on knowledge of the true, but unknown, clusters. Despite its conceptual appeal, the practical utility of these methods for post-clustering inference remains limited to extreme cases with extremely high signal vs. noise ratios, emphasizing the need for alternative methodologies that can navigate the complexities of unknown class structures more effectively.
@@
All codes and data needed to reproduce the results presented here are openly accessible from Zenodo with DOI 10.5281/zenodo.11207777 (Hivert et al., 2024b, ).
@@ 5 Acknowledgements
BH is supported partly by the Digital Public Health Graduate’s school, funded by the PIA $3$ (Investments for the Future – Project reference: 17-EURE-0019). The work was supported through the DESTRIER Inria Associate-Team from the Inria@SiliconValley program (analytical code: DRI-012215), and by the CARE project funded from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement no IMI2-101005077. The JU receives support from the European Union’s Horizon $2020$ research and innovation programme and EFPIA and Bill &\&& Melinda Gates Foundation, Global Health Drug Discovery Institute, University of Dundee. This work has received funding from the European Union’s Horizon $2020$ research and innovation programme under EHVA grant agreement no H2020-681032. This study was carried out in the framework of the University of Bordeaux’s France 2030 program / RRI PHDS. This work benefited from State aid managed by the Agence Nationale de la Recherche under the France 2030 program, reference AI4scMED ANR-22-PESN-0002. Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour.
Appendix S1 Independence proof of the Gaussian process
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\Sigma}_{p\times p}\right)$. Considering the fission process for Gaussian data described in Table $1$, we can decompose $\boldsymbol{X}$ into two new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ using a new random variable $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Sigma}\right)$. It follows form this decomposition that:
@@
$\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]$ $=\mathbb{E}\left[\left(\boldsymbol{X}+\tau\boldsymbol{Z}\right)\left(\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{Z}\right)^{t}\right]$
$=\mathbb{E}\left[\boldsymbol{X}\boldsymbol{X}^{t}-\frac{1}{\tau}\boldsymbol{X}\boldsymbol{Z}^{t}+\tau\boldsymbol{Z}\boldsymbol{X}^{t}-\boldsymbol{Z}\boldsymbol{Z}^{t}\right]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}\boldsymbol{Z}^{t}]+\tau\mathbb{E}[\boldsymbol{Z}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}]\mathbb{E}[\boldsymbol{Z}^{t}]+\tau\mathbb{E}[\boldsymbol{Z}]\mathbb{E}[\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]\qquad\mbox{since}\qquad\boldsymbol{X}\perp\!\!\!\perp\boldsymbol{Z}$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{Z}]=0$
@@
$\mathbb{V}\text{ar}(\boldsymbol{Z})=\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]-\mathbb{E}[\boldsymbol{Z}]\mathbb{E}[\boldsymbol{Z}^{t}]=\boldsymbol{\Sigma}$
$\iff\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]=\boldsymbol{\Sigma}+\mathbb{E}[\boldsymbol{Z}]\mathbb{E}[\boldsymbol{Z}^{t}]$
$\iff\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]=\boldsymbol{\Sigma}\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{Z}]=0$
@@
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{Z}\boldsymbol{Z}^{t}]-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=\boldsymbol{\Sigma}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}-\boldsymbol{\Sigma}-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=0$ (4)
Appendix S2 Impact of covariance estimation under the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$
Now, let suppose that we use $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\hat{\boldsymbol{\Sigma}}\right)$ to perform data fission. It follows from equation (4) that, since $\operatorname{\mathbb{V}ar}\left(\boldsymbol{Z}\right)=\hat{\boldsymbol{\Sigma}}$,
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\boldsymbol{\Sigma}-\hat{\boldsymbol{\Sigma}}$
Appendix S3 Derivation of $t$-test statistic under the null of no-cluster in the univariate data fission post-clustering setting

### v3

@@ 3.3 Application to single-cell RNA-seq data analysis
Single-cell RNA-seq (scRNA-seq) data analysis pipelines often involve an initial cluster analysis followed by differential analysis to identify marker genes and annotate cell populations based on gene expression. Given their overdispersed count nature, the negative binomial distribution is favored over the Gaussian distribution to model scRNA-seq data. Unfortunately, the negative binomial distribution also raises challenges for data thinning. The overdispersion parameter $\theta$ plays a role similar to the variance parameter in the Gaussian distribution, as it is considered known. Therefore, the quality of its estimation directly impacts $\operatorname{Cov}\left(X^{(1)},X^{(2)}\right)=\tau(1-\tau)\frac{\mu^{2}}{\theta}\left(1-\frac{\theta+1}{\widehat{\theta}+1}\right)$, where $\widehat{\theta}$ is an estimate of $\theta$ (Neufeld, Dharamshi, Gao & Witten 2024). Moreover, in negative binomial mixture models, conditional data thinning is also preferable to marginal thinning due to the component-specific nature of the overdispersion parameter (Li et al. 2018). However, applying conditional thinning requires accurate estimates of these parameters, which is challenging in practice since class memberships are unknown and must be inferred. This makes it difficult to ensure conditional independence and control Type I error in post-thinning analyses.
@@
We further illustrate the necessity of applying conditional data thinning (with the associated conditional overdispersion) to ensure Type I error control in the post-clustering inference setting with numerical simulations. We generated $n=100$ observations from a two-component negative binomial mixture: $0.5\text{NegBin}(\mu_{1},\theta_{1})+0.5\text{NegBin}(\mu_{2},\theta_{2})$ with component parameters $(\mu_{1},\theta_{1})=(5,5)$ and $(\mu_{2},\theta_{2})=(60,40)$. We conducted similar post-clustering inference as in Figure 1. In Figure 3A, applying the $k$-means with $G=3$ clusters reveals that the first mixture component is erroneously split into 2 clusters (labeled $C_{1}$ and $C_{3}$). We then assessed the Type I error rate associated with the Wilcoxon test between these spurious clusters when applying both marginal and conditional data thinning. First, we applied conditional data thinning using oracle estimates $\widetilde{\theta}_{g}$, $g=1,2$, representing true intra-component overdispersion parameters (infeasible in real-life application where the true class memberships are unknown). We compared those results with marginal data thinning with its associated marginal $\widehat{\theta}$. Of note, all the overdispersion estimations were performed using Maximum Likelihood Estimation. Figure 3B presents the QQ-plot against the Uniform distribution of the associated Wilcoxon p-values across 1,0001,\!000 simulations. Similarly to the Gaussian setting, achieving uniformly distributed p-values is only possible through conditional data thinning using oracle conditional overdispersion estimates $\widetilde{\theta}_{g}$. Marginal negative binomial data thinning does not guarantee conditional independence between $X^{(1)}$ and $X^{(2)}$, leading to inflated Type I error rates in post-clustering inference – particularly when a true component is erroneously split into multiple clusters.
@@
Figure 3: Challenges in estimating overdispersion, a gene-specific parameter, for negative binomial data thinning. A: spurious clustering in simulated data. B: $Q$-plot of Wilcoxon p-values when testing between two spurious clusters. C: concordance between gene-wise overdispersions estimated within 4 different cell populations and in neutrophils.
@@
Extending our investigation from simulated scenarios to real-life applications, we used a single-cell RNA-seq dataset from the Tabula Sapiens Consortium (The Tabula Sapiens Consortium* 2022) to explore the practical challenges of estimating overdispersion. Our analysis focused on five distinct cell populations from bone marrow: 2,5602,\!560 neutrophils, $105$ macrophages, $386$ monocytes, $454$ granulocytes, and $833$ CD4 T cells – all collected from a single donor (donor TSP14). In this controlled setting, where cell types were known, we succeeded in estimating the overdispersion of 8,3338,\!333 genes for each cell type using the variance stabilizing transformation implemented in the sctransform package (Choudhary & Satija 2022). Figure 3C illustrates the comparison of gene overdispersion when estimated solely in neutrophils versus the overdispersion estimated for the same genes in the other four cellular populations. Root Mean Squared Error (RMSE) values were computed to quantify the agreement between estimations. Our findings reveal that overdispersion is specific to each cell population, as evidenced by a notable deviation from the diagonal and relatively high RMSE values. This highlights the need for conditional data thinning in such an application. Combined with our simulation studies, these results demonstrate how Type I error can easily be inflated in real-life applications of data thinning for scRNA-seq data analysis, particularly due to the difficulty in providing an unbiased estimate of gene overdispersion to be used in conditional data thinning.
@@ 4 Discussion
We highlight here the practical limitations inherent in data fission and its extension, data thinning, for post-clustering inference challenges. An important issue is the assumption of a homogeneous data distribution, which implies the absence of true clusters in the data. To address this limitation and adapt to scenarios with true classes, a shift towards mixture models becomes imperative. However, these models lack a predefined decomposition through data fission or thinning. We have shown that applying data fission or thinning without accounting for the mixture structure – that is, using marginal approaches – leads to inflated Type I error rates and quite low statistical power in post-clustering inference.
@@
We have proposed conditional data fission and data thinning, which perform the decomposition at the component level, and demonstrated their theoretical validity. However, theses approaches rely on prior knowledge of the mixture components scale parameters, such as variances in the Gaussian distribution or the overdispersion in the negative binomial distribution. However in real-life applications, these parameters are unknown. Adequately estimating these parameters becomes intricate in the presence of true clusters, given their component-specific nature; meanwhile the quality of the estimation of those parameters is directly linked with the covariance between the new random variables, $X^{(1)}$ and $X^{(2)}$, decomposing the original data. Only unbiased estimation of these parameters ensures the conditional independence between $X^{(1)}$ and $X^{(2)}$. That independence is paramount for post-clustering inference to adequately control the Type I error rate. Thus, although conditional data fission offers a promising solution for valid post-clustering inference, it creates a circular dependency. The clusters must be known in advance to perform the decomposition, even though this decomposition is intended to enable valid inference about such estimated clusters (see Supplementary Figure 1).
@@
In the Gaussian framework, we theoretically quantify the relationship between the relative bias in variance estimation and the associated Type I error in post-clustering $t$-tests. In practice, our simulation results suggest that a small relative bias can be acceptable while still achieving effective Type I error control. These first results pave the way for defining a principled approach to tuning the hyperparmater $\tau$ in data fission and data thinning for post-clustering inference to optimize statistical power.
@@
Finally, we demonstrate that the results derived from the Gaussian distribution context are readily applicable to the negative binomial distribution commonly employed in modeling RNA-seq data. Specifically, we illustrate on real data that overdispersion is also component-specific. Therefore, without knowing the true class memberships, data thinning cannot ensure the needed independence between clustering and differential testing.
To avoid the need for prior knowledge of component-specific variances in Gaussian mixture settings, one may consider adopting a heteroscedastic model with observation-level means and variances. Within this framework, variances can be approximated using plug-in estimators, such as non-parametric estimates of local variability. This approach is more consistent with the distributional assumptions underlying data fission and thinning. Moreover, such heteroscedastic models naturally encompass mixture models as a special case, where the local variance reflects the component-specific structure. Thus, they might hold a path for approximating conditional data fission, without requiring explicit knowledge of the mixture parameters. Nevertheless, their practical deployment presents notable challenges. Chief among them is the sensitivity of non-parametric variance estimates to the choice of bandwidth. Ideally, the bandwidth should be small enough to capture only observations from the same component. Yet, this implicitly assumes knowledge of the true mixture structure, and its effectiveness ultimately depends on the degree of separation between components.
@@
In practice, the application of data fission or data thinning for post-clustering inference appears to be akin to a self-referential loop generating circular reasoning. While introduced as a solution for addressing challenges in post-clustering inference, all strategies that could theoretically ensure independence between the two stages of the analysis ultimately rely on knowledge of the true, but unknown, class memberships. Despite its conceptual appeal, the practical utility of these methods for post-clustering inference remains limited to extreme cases with extremely high signal vs. noise ratios, emphasizing the need for alternative methodologies that can navigate the complexities of unknown class structures more effectively.
@@
All codes and data needed to reproduce the results presented here are openly accessible from Zenodo with DOI 10.5281/zenodo.11207776.
@@ 5 Acknowledgements
BH is supported partly by the Digital Public Health Graduate’s school, funded by the PIA $3$ (Investments for the Future – Project reference: 17-EURE-0019). The work was supported through the DESTRIER Inria Associate-Team from the Inria@SiliconValley program (analytical code: DRI-012215), and by the CARE project funded from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement no IMI2-101005077. The JU receives support from the European Union’s Horizon $2020$ research and innovation programme and EFPIA and Bill &\& Melinda Gates Foundation, Global Health Drug Discovery Institute, University of Dundee. This work has received funding from the European Union’s Horizon $2020$ research and innovation programme under EHVA grant agreement no H2020-681032. This study was carried out in the framework of the University of Bordeaux’s France 2030 program / RRI PHDS. This work benefited from State aid managed by the Agence Nationale de la Recherche under the France 2030 program, reference AI4scMED ANR-22-PESN-0002. Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour. The authors disclose the use of generative AI (ChatGPT, OpenAI, 2025) to help refine the phrasing of some sections of the manuscript.
Supplementary Materials
Appendix 1 Independence proof of the Gaussian process
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\Sigma}_{p\times p}\right)$. Considering the fission process for Gaussian data described in Table $1$, we can decompose $\boldsymbol{X}$ into two new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ using a new random variable $\boldsymbol{W}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Sigma}\right)$. It follows form this decomposition that:
@@ Appendix 1 Independence proof of the Gaussian process
$\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]$ $=\mathbb{E}\left[\left(\boldsymbol{X}+\tau\boldsymbol{W}\right)\left(\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{W}\right)^{t}\right]$
$=\mathbb{E}\left[\boldsymbol{X}\boldsymbol{X}^{t}-\frac{1}{\tau}\boldsymbol{X}\boldsymbol{W}^{t}+\tau\boldsymbol{W}\boldsymbol{X}^{t}-\boldsymbol{W}\boldsymbol{W}^{t}\right]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}\boldsymbol{W}^{t}]+\tau\mathbb{E}[\boldsymbol{W}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}]\mathbb{E}[\boldsymbol{W}^{t}]+\tau\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]\qquad\mbox{since}\qquad\boldsymbol{X}\perp\!\!\!\perp\boldsymbol{W}$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{W}]=0$
@@
$\mathbb{V}\text{ar}(\boldsymbol{W})=\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]-\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}$
$\iff\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}+\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{W}^{t}]$
$\iff\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{W}]=0$
@@
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=\boldsymbol{\Sigma}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}-\boldsymbol{\Sigma}-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=0$ (4)
Appendix 2 Covariance Structures in Conditional and Marginal Gaussian Data Fission
In the following, we assume that $\boldsymbol{X}_{i}$ is distributed according to a Gaussian mixture with conditional variance $\boldsymbol{\Sigma}_{g}$ and marginal variance $\boldsymbol{\Sigma}$.
2.1 Conditional Data Fission
Conditional data fission guarantees conditional independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ due to the properties of Gaussian decomposition (see Section $1$ of Supplementary Materials). However, marginal dependence may still arise, as shown below:
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)$ $=\mathbb{E}\left[\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}\right)\right]+\operatorname{Cov}\left(\mathbb{E}\left[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}\right],\mathbb{E}\left[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}\right]\right)$
$=\operatorname{Cov}\left(\mathbb{E}[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}],\mathbb{E}[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}]\right)\quad\text{(by conditional independence)}$
$=\operatorname{Cov}(\boldsymbol{\mu}_{Z},\boldsymbol{\mu}_{Z})$
$=\mathbb{V}\text{ar}(\boldsymbol{\mu}_{Z}),$
where $\boldsymbol{\mu}_{Z}=\mathbb{E}[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}]=\mathbb{E}[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}]$. Since $Z_{i}$ is discrete, we have $\boldsymbol{\mu}_{Z}=\boldsymbol{\mu}_{g}$ with probability $\pi_{g}$, so:
$\mathbb{E}[\boldsymbol{\mu}_{Z}]=\sum_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g},\quad\text{and}\quad\mathbb{V}\text{ar}(\boldsymbol{\mu}_{Z})=\sum_{g=1}^{G}\pi_{g}(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])^{T}.$
Therefore, the marginal covariance between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ in conditional fission equals the variance of the component means:
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)=\sum_{g=1}^{G}\pi_{g}(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])^{T}.$
2.2 Marginal Data Fission
Marginal covariance
We now demonstrate that marginal data fission induces no marginal correlations between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$. Let:
$\boldsymbol{X}_{i}^{(1)}=\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\quad\boldsymbol{X}_{i}^{(2)}=\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i},$
where $\boldsymbol{X}_{i}$ has covariance matrix $\boldsymbol{\Sigma}$ and $\boldsymbol{W}_{i}\overset{\text{iid}}{\sim}\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$, with $\boldsymbol{X}_{i}\perp\!\!\!\perp\boldsymbol{W}_{i}$.
$\operatorname{Cov}(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)})$ $=\operatorname{Cov}(\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\boldsymbol{X}_{i}-\tfrac{1}{\tau}\boldsymbol{W}_{i})$
$=\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{X}_{i})+\tau\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{X}_{i})-\tfrac{1}{\tau}\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{W}_{i})-\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{W}_{i})$
$=\boldsymbol{\Sigma}-\boldsymbol{\Sigma}$
$=\mathbf{0}.$
Conditional covariance
Next, we quantify the conditional dependence given the class membership $Z_{i}=g$:
$\operatorname{Cov}(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g)$ $=\operatorname{Cov}(\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\boldsymbol{X}_{i}-\tfrac{1}{\tau}\boldsymbol{W}_{i}\mid Z_{i}=g)$
$=\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{X}_{i}\mid Z_{i}=g)+\tau\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{X}_{i}\mid Z_{i}=g)$
$\quad-\tfrac{1}{\tau}\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{W}_{i}\mid Z_{i}=g)-\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{W}_{i}\mid Z_{i}=g)$
$=\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma},$
where $\boldsymbol{\Sigma}_{g}$ denotes the class-specific covariance. Thus, unless all components share the same covariance structure ($\boldsymbol{\Sigma}_{g}=\boldsymbol{\Sigma}$), marginal data fission introduces conditional correlation despite no marginal dependence.
Appendix 3 Impact of covariance estimation under the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$